# Parallelizing the Approximate Minimum Degree Ordering Algorithm: Strategies and Evaluation

### v1

@@ 3.3.1. Concurrent Connection Updates
3.3.2. Concurrent Approximate Degree Lists
In the sequential AMD implementation from SuiteSparse (Davis, 2019), pivot selection is facilitated by maintaining $n$ degree lists. Each list corresponds to a specific approximate degree and stores the associated variables using a doubly linked list. When a pivot is selected, it is removed from the degree list, and for each variable in its neighborhood, it is moved from its current list to the new one based on its updated approximate degree. No dynamic memory allocation is required, as the total number of entries across all lists is bounded by $n$. However, this data structure is not amenable to efficient parallelization due to high contention when multiple threads attempt to access the same doubly linked list.
Instead, we let each thread $tid$ maintain its own approximate degree lists $dlist_{tid}(0,\cdots,n-1)$, along with a shared $affinity$ array to track which thread holds the most up-to-date information for each variable, and a private $loc$ array to record local degree list location of each variable. Both the $affinity$ and $loc$ arrays are initialized to $-1$. The use of a distance-2 independent set of pivots ensures that each variable is associated with at most one pivot—and thus with a single thread—during any given elimination step. As shown in Algorithm 2, when a thread attempts to remove a variable $v$, it sets $affinity(v)$ to $-1$, effectively invalidating all entries of $v$ across all threads without physically modifying the lists. When inserting $v$, the thread first checks if $v$ is currently used in its own lists based on $loc(v)$. If so, the stale entry of $v$ is explicitly removed to reclaim memory. In all cases, the variable $v$ is inserted into the appropriate list for the thread, with $loc(v)$ updated accordingly and $affinity(v)$ set to $tid$ to reflect the change. Other invalid entries, detected via mismatched affinity, are lazily reclaimed during list traversal for pivot selection. This design avoids contention for degree list maintenance, with the only necessary coordination occurring when determining the minimum approximate degree from the per-thread lists at the start of each elimination step.
3.4. Parallel Distance-2 Independent Set
To parallelize the selection of a distance-2 independent set for multiple elimination, we first collect candidate variables whose approximate degrees fall within the relaxed threshold described in Section 3.2. This collection is performed in parallel from each thread’s degree lists. To limit overhead compared to the sequential AMD algorithm, we restrict each thread to collect at most $lim$ candidates. The impact of this parameter will be analyzed in Section 4.5.
To select a distance-2 independent set, we perform a single iteration of the distance-2 analog of Luby’s algorithm (Luby, 1985) as shown in Algorithm 3. For each candidate gathered from the approximate degree lists, a random number is generated and compared against the candidates within its distance-2 neighborhood, with ties broken by candidate indices. A candidate is considered valid if and only if it has the smallest value among the candidates within its distance-2 neighborhood. To enable parallelism, Algorithm 3 includes appropriate atomic operations and synchronization barriers. We intentionally avoid running the full algorithm because a maximal distance-2 independent set is unnecessary; as long as the set obtained in the first iteration is sufficiently large, it provides enough parallelism for the core AMD algorithm.
3.5. Overall Parallel AMD Algorithm
Algorithm 4 summarizes our strategies used to parallelize the AMD algorithm. These techniques collectively enable an efficient and scalable implementation, addressing long-standing challenges in parallelizing the AMD algorithm.
Algorithm 4 Parallel Approximate Minimum Degree Algorithm
$|V|>0$
@@ 3.5. Overall Parallel AMD Algorithm
3:     Eliminate $p$ using the core AMD algorithm
4:      with concurrent connection updates and
5:      concurrent approximate degree lists
6:      (see Section 3.3)
@@
4. Experiments
4.1. Evaluation Platforms
We evaluated the performance of our parallel approximate minimum degree algorithm on a CPU node of NERSC Perlmutter, an HPE Cray EX supercomputer. One CPU node has 128 physical cores divided between two AMD EPYC 7763 (Milan) CPUs with 512 GB of DDR4 memory in total. We used a single CPU with 64 cores in our evaluation. Our implementation was written in C++ using OpenMP for shared memory parallelism and compiled with GCC 13.2.1 using the flags -O3 -fopenmp, and it is available online at https://github.com/PASSIONLab/ParAMD. As a baseline, we used the sequential implementation from SuiteSparse (v7.10.1) (Davis, 2019). To the best of our knowledge, no existing parallel implementation of the approximate minimum degree algorithm (AMD) is available for comparison.
4.2. Matrix Suite
The sparse matrices used in our experiments were selected from the SuiteSparse Matrix Collection (Davis and Hu, 2011) to represent a diverse range of large-scale linear systems arising from real-world applications. These matrices are summarized in Table 4. Although some matrices are nonsymmetric, the AMD algorithm can still be applied by considering the sum of the matrix and its transpose, as is done in the SuiteSparse AMD implementation. In fact, this pre-processing step is performed in SuiteSparse AMD regardless of whether the input matrix is symmetric or not. For fairness in end-to-end comparison, we also parallelize this pre-processing step using simple atomic operations and include its runtime in the reported results.
Table 4. Matrices used in our experiments sorted by #nonzeros. Matrices pre2, HV15R, and stokes are nonsymmetric.
Matrix Name #rows #nonzeros Problem Type
pre2 659K 5.83M circuit simulation
nd24k 72.0K 28.7M 3D mesh
ldoor 952K 42.5M structural
dielFilterV3real 1.10M 89.3M electromagnetics
Flan_1565 1.56M 114M structural
HV15R 2.02M 283M fluid dynamics
Queen_4147 4.15M 317M structural
stokes 11.4M 349M semiconductor
nlpkkt240 28.0M 761M optimization
4.3. End-to-End Comparison
We now present the performance of our parallel AMD implementation compared to the SuiteSparse baseline, as summarized in Table 5. Matrices were randomly permuted five times with all methods evaluated on the same set of permutations to decouple tie-breaking issues explained in Section 2.5.4. We used the default relaxation factor $mult=1.1$ and set the limitation factor $lim$ to $8192$ divided by the number of threads, following the heuristic that this configuration targets a fill-in ratio of $\times1$ over the SuiteSparse baseline while ensuring a sufficiently large pool of candidates is considered for the distance-2 independent set in each elimination step. Note that with this choice of $lim$, the number of candidates remains nearly constant, regardless of the number of threads used.
Table 5. End-to-end comparison between SuiteSparse AMD and our parallel AMD implementation with 1 thread (1t) and 64 threads (64t). Matrices were randomly permuted five times with all methods evaluated on the same set of permutations to decouple tie-breaking issues. We report the median ordering time and fill-in across the five runs. Speedups and fill-in ratios are computed directly based on the medians.
Matrix Name Ordering Time (sec) Our Speedup #Fill-ins Our Fill-in Ratio
SuiteSparse Ours over SuiteSparse SuiteSparse Ours over SuiteSparse
1t 64t 1t 64t 1t 64t 1t 64t
pre2 2.01 4.70 0.62 0.43$\times$ 3.25$\times$ 1.84e+08 1.93e+08 1.91e+08 1.05$\times$ 1.04$\times$
nd24k 0.82 1.01 0.26 0.81$\times$ 3.15$\times$ 5.09e+08 5.23e+08 5.07e+08 1.03$\times$ 1.00$\times$
ldoor 1.37 1.35 0.29 1.01$\times$ 4.79$\times$ 1.52e+08 1.57e+08 1.58e+08 1.03$\times$ 1.04$\times$
dielFilterV3real 2.63 2.78 0.46 0.95$\times$ 5.74$\times$ 9.65e+08 1.06e+09 1.07e+09 1.10$\times$ 1.11$\times$
Flan_1565 4.20 4.51 0.80 0.93$\times$ 5.28$\times$ 3.72e+09 3.96e+09 4.01e+09 1.07$\times$ 1.08$\times$
HV15R 11.75 13.86 2.15 0.85$\times$ 5.46$\times$ 5.65e+10 5.65e+10 5.80e+10 1.00$\times$ 1.03$\times$
Queen_4147 13.82 17.16 2.46 0.81$\times$ 5.61$\times$ 7.72e+10 8.35e+10 8.63e+10 1.08$\times$ 1.12$\times$
stokes 79.32 121.06 10.63 0.66$\times$ 7.46$\times$ 7.76e+10 8.23e+10 8.00e+10 1.06$\times$ 1.03$\times$
nlpkkt240 296.67 433.95 35.75 0.68$\times$ 8.30$\times$ 1.60e+12 1.83e+12 1.79e+12 1.14$\times$ 1.12$\times$
Table 5 shows that, when running with a single thread, our implementation incurs overhead, resulting in slower ordering time. This is primarily due to the inclusion of Algorithm 3 for selecting distance-2 independent sets, a step not required in the original AMD algorithm. However, it is this modification that enables effective parallelization, achieving up to an 8.30×8.30$\times$8.30 × speedup over SuiteSparse and up to a 12.21×12.21$\times$12.21 × self-speedup when using 64 threads. Our fill-in ratio is consistently around $\times1$ over SuiteSparse, reflecting the effect of the relaxation factor $mult$ in avoiding the selection of overly suboptimal pivots.
4.4. Runtime Breakdown and Parallelism Analysis
To gain deeper insights into the performance of our parallel AMD implementation, we present a runtime breakdown in Figure 5 using the default parameters.
First, we observe that the pre-processing step, which computes $A+A^{T}$, becomes a bottleneck for some matrices and does not scale well with increasing thread count. This behavior is expected due to high contention and irregular memory access patterns inherent to the operation. However, this step is unnecessary for symmetric matrices and, for nonsymmetric matrices including pre2, HV15R, and stokes, its impact is minimal. We include it in the runtime solely for a fair end-to-end comparison with the SuiteSparse baseline.
Second, while the core of the AMD algorithm is highly parallelizable in theory, it does not achieve perfect scaling in practice. For example, on the nd24k matrix, we observe only a 3.7×3.7$\times$3.7 × speedup when using 64 threads for the core AMD part. This limited scalability is explained by the distribution of the sizes of the distance-2 independent sets, as shown in Figure 6, which directly affects the amount of parallel work available in each elimination step. Unsurprisingly, the distance-2 independent sets selected for nd24k are relatively small compared to other matrices, resulting in underutilization of threads. Moreover, across all matrices, a significant portion of the distance-2 independent sets still have sizes less than 64, further illustrating why the core AMD algorithm does not scale perfectly despite being embarrassingly parallel. As a future direction, we propose dynamically adapting the relaxation and limitation factors during elimination, particularly when low workload is detected, to extract greater parallelism when needed.
Figure 5. Runtime breakdown of our parallel AMD algorithm as the number of threads scales from 1 to 64.
@@ 4.4. Runtime Breakdown and Parallelism Analysis
Runtime breakdown of our parallel AMD algorithm as the number of threads scales from 1 to 64.
Figure 6. Violin plots showing the distribution of the sizes of distance-2 independent sets across elimination steps. The plots are cropped to reflect the actual range of the data. The dotted line marks the threshold of 64, representing the minimum size needed to fully utilize all 64 threads.
@@
Violin plots showing the distribution of distance-2 independent set sizes across elimination steps. The dotted line marks the threshold of 64, representing the minimum size needed to fully utilize all 64 threads.
Figure 7. Impact of the relaxation factor $mult$ and the limitation factor $lim$ on core AMD runtime, distance-2 independent set selection time, and ordering quality. Results are shown for matrices nd24k and nlpkkt240, representing the worst and best scalability cases, respectively. All experiments were conducted using 64 threads. Lighter colors indicate better performance.
@@
Impact of the relaxation factor $mult$ and limitation factor $lim$ on core AMD runtime, distance-2 independent set selection time, and ordering quality. Results are shown for matrices nd24k and nlpkkt240, representing the worst and best scalability cases, respectively. All experiments were conducted using 64 threads. Lighter colors indicate better performance.
4.5. Relaxation and Limitation Factors
@@
We now evaluate the impact of the relaxation factor $mult$ and the limitation factor $lim$. Figure 7 illustrates how these parameters affect the runtime of the core AMD algorithm, the runtime of distance-2 independent set selection, and the overall ordering quality. Our analysis focuses on two representative matrices using 64 threads: nd24k, which demonstrates the worst scalability, and nlpkkt240, which shows the best scalability, in Table 5.
@@
The impact of the relaxation factor $mult$ is more pronounced and consistent. When $mult$ is too small, the algorithm lacks sufficient parallelism, leading to poor performance. On the other hand, excessively large values of $mult$ degrade the ordering quality due to the inclusion of suboptimal pivots.
@@
In contrast, the impact of the limitation factor $lim$ is less straightforward. Its effects are more nuanced and do not follow a universal trend. This is because adjusting $lim$ affects several aspects of the algorithm, including but not limited to the total number of elimination steps, opportunities to apply other orthogonal techniques discussed in Section 2.4, and the amount of computation available to hide memory access latency. As these factors impact different matrices to varying degrees, the limitation factor $lim$ produces complex behavior with no single dominant pattern.
@@
Nevertheless, Figure 7 suggests that the optimal configuration occurs around $mult=1.2$ and $lim=128$. After further fine-grained tuning across our dataset, we found that setting $mult=1.1$ provides a better balance between runtime and ordering quality. To ensure that the number of candidates for the distance-2 independent set remains consistent regardless of the number of threads, we finalize $lim$ as 8192 divided by the number of threads. This results in $lim=128$ when using 64 threads, aligning with the identified optimal value. These together result in the default settings we used for the other experiments.
5. Conclusion and Future Work
@@
We have presented the first scalable shared memory AMD implementation. By incorporating distance-2 independent sets into multiple elimination and designing specialized concurrent data structures tailored to the AMD algorithm, we have enabled effective parallelism while reducing contention. Our experiments have demonstrated that this approach scales and consistently outperforms the long-standing AMD implementation from SuiteSparse, with only modest and controllable overhead in fill-in.

### v2

@@ Abstract
[…]
4.4 Runtime Breakdown and Parallelism Analysis.
To gain deeper insights into our parallel AMD implementation, we present a runtime breakdown in Figure 4.1 using the default parameters.
First, when running with a single thread, our implementation is slower than the baseline. This is primarily due to the inclusion of Algorithm 2 for selecting distance-2 independent sets, a step not required in the original AMD algorithm. However, it is this modification that enables effective parallelization that makes our implementation scalable.
Second, we observe that the pre-processing step, which computes $|A|+|A^{T}|$, becomes a bottleneck for some matrices and does not scale well with increasing thread count. This behavior is expected due to high contention and irregular memory access inherent to the operation. Therefore, we suggest users skip this pre-processing step if the input is known to be symmetric.
Third, while the core AMD part is highly parallelizable in theory, it does not achieve perfect scaling in practice. This limited scalability is explained by the distribution of the sizes of the distance-2 independent sets, as shown in Figure 4.2. Notably, for the matrix that scales the worst, nd24k, the distance-2 independent sets selected are relatively small compared to other matrices, resulting in underutilization of threads. Moreover, across all matrices, a significant portion of the distance-2 independent sets still have sizes less than 64, further illustrating why the core AMD algorithm does not scale perfectly despite being embarrassingly parallel.
Figure 4.1: Runtime breakdown of our parallel AMD algorithm as the number of threads scales from 1 to 64.
Figure 4.2: Violin plots showing the distribution of the sizes of distance-2 independent sets across elimination steps. The plots are cropped to reflect the actual range of the data. The dotted line marks the threshold of 64, representing the minimum size needed to fully utilize all 64 threads.
@@
Figure 4.3: Impact of the relaxation factor multmult and the limitation factor limlim. Results are shown for matrices nd24k and nlpkkt240, representing the worst and best scalability cases, respectively. All experiments were conducted using 64 threads. Lighter colors indicate better performance.
4.5 Relaxation and Limitation Factors.
@@
We now evaluate the impact of the relaxation factor $\mathit{mult}$ and the limitation factor $\mathit{lim}$. Figure 4.3 shows how they affect the core AMD runtime, the runtime of distance-2 independent set selection, and the ordering quality. Our analysis focuses on two representative matrices using 64 threads: nd24k and nlpkkt240, which demonstrates the worst and the best scalability, respectively.
@@
The impact of the relaxation factor $\mathit{mult}$ is more pronounced and consistent. When $\mathit{mult}$ is too small, the algorithm lacks sufficient parallelism, leading to poor performance. On the other hand, excessively large values of $\mathit{mult}$ degrade the ordering quality due to the inclusion of suboptimal pivots.
@@
In contrast, the impact of the limitation factor $\mathit{lim}$ is less straightforward. This is because adjusting $\mathit{lim}$ affects several aspects of the algorithm, including the number of elimination steps, opportunities to apply other techniques discussed in Section 2.4, and the amount of computation available to hide memory access latency. As these factors impact different matrices to varying degrees, the limitation factor $\mathit{lim}$ produces complex behavior with no single dominant pattern.
@@
Nevertheless, Figure 4.3 suggests that the optimal configuration occurs around $\mathit{mult}=1.2$ and $\mathit{lim}=128$. After further fine-grained tuning across our dataset, we found that setting $\mathit{mult}=1.1$ provides a better balance between runtime and ordering quality. To ensure that the number of candidates for the distance-2 independent set remains constant regardless of the number of threads, we finalize $\mathit{lim}$ as 8192 divided by the number of threads. This results in $\mathit{lim}=128$ when using 64 threads, aligning with the identified optimal value. These together result in the default settings.
Table 4.3: End-to-end comparison of ordering methods for solving sparse symmetric positive definite systems. Each matrix is randomly permuted five times, with all methods evaluated on the same set of permutations to avoid the tie-breaking issues discussed in Section 2.5.4. All orderings are computed on the CPU—SuiteSparse AMD uses single thread and the others use 64 threads—while the reordered systems are solved on the GPU using cuDSS in double precision.
Matrix SuiteSparse AMD Ours cuDSS ND
Ordering GPU Solver Ordering GPU Solver Ordering GPU Solver
Time (sec) Time (sec) Time (sec) Time (sec) Time (sec) Time (sec)
nd24k 0.82 $\pm$ 0.00 1.97 $\pm$ 0.07 0.26 $\pm$ 0.00 1.97 $\pm$ 0.09 2.38 $\pm$ 0.13 1.27 $\pm$ 0.02
ldoor 1.42 $\pm$ 0.05 3.03 $\pm$ 0.01 0.32 $\pm$ 0.03 3.07 $\pm$ 0.01 0.67 $\pm$ 0.01 2.48 $\pm$ 0.01
Flan_1565 4.85 $\pm$ 0.56 18.92 $\pm$ 0.22 0.95 $\pm$ 0.27 19.62 $\pm$ 0.23 2.41 $\pm$ 0.01 9.21 $\pm$ 0.02
Cube5317k 13.94 $\pm$ 0.92 43.90 $\pm$ 0.37 2.90 $\pm$ 0.67 44.33 $\pm$ 0.41 7.43 $\pm$ 0.04 33.45 $\pm$ 0.21
4.6 End-to-End Solver Performance.
Lastly, we evaluated end-to-end solver performance by solving the reordered systems using orderings computed by SuiteSparse AMD, our parallel AMD implementation, and a multi-threaded nested dissection (ND) method. For the solver, we focused on symmetric positive definite systems—eliminating the need for numerical pivoting—and employed cuDSS [cudss] (v0.7.1) on an NVIDIA A100 80GB GPU in double precision. The multi-threaded ND used in our evaluation is the one that comes with cuDSS, which is a customized variant of METIS [karypis1997metis]. Note that all orderings are computed on the CPU—SuiteSparse AMD uses single thread and the others use 64 threads—while the reordered systems are solved on the GPU. For this experiment, we restricted our evaluation to only four matrices in our dataset as the input matrices need to be symmetric positive definite and their Cholesky factors need to fit within the available GPU memory.
From Table 4.3, we observe that the end-to-end performance improves with our parallel AMD compared to SuiteSparse AMD. While GPU running time increases slightly due to extra fill-ins, this overhead is outweighed by the reduced ordering time. Although the extra fill-ins are generally not problematic, we acknowledge that they may cause out-of-memory issues in some cases. Taking the Serena matrix as an example, its Cholesky factor barely fits on the GPU when using SuiteSparse AMD, but exceeds the limit when using our parallel AMD. Therefore, it is omitted here. On the other hand, our parallel AMD outperforms multi-threaded ND for the reordering phase, with the slowdown occurring in the GPU solver time. Systems reordered by multi-threaded ND can be solved faster because they have fewer fill-ins, as summarized in Table 4.4. Moreover, the solver benefits from the partitions produced by ND, which enable better parallelism during factorization. We believe a hybrid approach combining our method with multi-threaded ND could yield a new parallel ordering that further enhances end-to-end solver performance.
Table 4.4: Comparison of the #Fill-ins by using the orderings from SuiteSparse AMD, our parallel AMD, and cuDSS ND.
Matrix #Fill-ins
SuiteSparse Ours cuDSS
AMD ND
nd24k 5.03e+08 5.18e+08 3.21e+08
ldoor 1.52e+08 1.57e+08 1.41e+08
Flan_1565 3.71e+09 3.94e+09 1.46e+09
Cube5317k 6.29e+09 6.54e+09 4.02e+09
5 Conclusion and Future Work.
@@
We have presented the first scalable shared memory AMD implementation. By incorporating distance-2 independent sets into multiple elimination and designing specialized concurrent data structures tailored to the AMD algorithm, we have enabled effective parallelism and reduced contention. Our experiments have demonstrated that this approach scales and consistently outperforms the AMD implementation from SuiteSparse, with only modest and controllable overhead in fill-in.